# Alignment-Sensitive Minimax Rates for Spectral Algorithms with Learned Kernels

### v2

@@ C.2 Generalized signal results by dynamic equation analysis
For the choice $b_{0}=cD^{\frac{D+1}{D+2}}\varepsilon^{\frac{1}{D+2}}$, the second term on the right-hand side of Equation˜48 is dominated by the right-hand side of Equation˜49, and we can choose $c$ small enough so that the summation of the two terms is bounded by $b_{0}^{-D}\varepsilon^{-1}\ln n$. This justifies our choice of $t(\varepsilon)$.
@@
Since the events in Lemma C.9 hold, we have $|\xi_{j}|\leq 2(C_{\text{proxy}})^{-1/2}\sqrt{\frac{\ln\tilde{j}+\ln n}{n}}$, where $\tilde{j}=\frac{\tilde{d}}{\lambda_{j}}$. Since $j\in R$, we have $|\theta_{j}^{*}|\leq\tilde{\sigma}\leq\sqrt{c^{\prime}}\varepsilon$. Since $b_{0}^{D}t(\varepsilon)=\varepsilon^{-1}\ln n$, we can check that $t$ is no more than the hitting time $T_{2}$ defined by Equation˜67 in Lemma˜C.8 as follows.
@@ C.3 Conservation quantity
We omit the subscript $j$ in the following two sections C.3 and C.4 because all the proofs are similar for $j=1,2,\cdots,d$. By Equation˜38, it is easy to see that
@@ C.4 Multi-layer dynamic
Consider the equation Equation˜38. Suppose $y>0$.
@@
Items 1 and 2 are directly implied from Equation˜38. Item 3 is implied by Item 2.
To prove Item 4, consider Equation˜52, from which we have
@@
Consider the equation Equation˜38. Suppose $\theta^{*}\geq 8\kappa$ (similar results hold for $\theta^{*}\leq-8\kappa$ by symmetry). Suppose $t_{0}\geq 0$ such that $0\leq\theta\left(t_{0}\right)<\frac{1}{4}\theta^{*}$. Define
@@
From Equation˜38, we have
@@
Case 1: If $a_{0}<b_{0}/\sqrt{D}$, $r=a_{0}$ and $R=b_{0}/\sqrt{D}$. Note $a\geq\beta$ in Equation˜51. We use Equation˜56 to get
@@
Case 2: If $a_{0}>b_{0}/\sqrt{D}$, $R=a_{0}$ and $r=b_{0}/\sqrt{D}$. We use $b\geq\sqrt{D}\beta$ in Equation˜51 together with $a>a_{0}$, $b>\sqrt{D}|\beta|$ and Equation˜56 to get that
@@
Stage 3: If $T^{\textrm{sig}}\leq T^{\textrm{pos},2}$, then we can use the for $T^{\textrm{sig}}$ in Stage 2 as a bound for $T^{\textrm{sig}}$. Now, we consider the case $T^{\textrm{pos},2}<T^{\textrm{sig}}$. We combine Equation˜51 with $a>|\beta|$, $b>\sqrt{D}|\beta|$, and Equation˜56 to get
@@
We now bound $T^{\mathrm{\textrm{sig}}}$ using the summation of Equation˜57, Equation˜60, and Equation˜58 if $a_{0}<b_{0}/\sqrt{D}$, or the summation of Equation˜57, Equation˜60, and Equation˜59 if $a_{0}>b_{0}/\sqrt{D}$.
If $a_{0}<b_{0}/\sqrt{D}$, we can bound the right hand sides of Equation˜57 and Equation˜60 by $8(\theta^{*})^{-1}b_{0}^{-D}$ and $4(\theta^{*})^{-1}b_{0}^{-D}$ respectively.
If $a_{0}>b_{0}/\sqrt{D}$, we can bound the right hand sides of Equation˜57 and Equation˜60 by $8(\theta^{*})^{-1}D^{-\frac{1}{2}}a_{0}^{-1}b_{0}^{-D+1}$ and $4(\theta^{*})^{-1}D^{-\frac{1}{2}}a_{0}^{-1}b_{0}^{-D+1}$ respectively. Furthermore, if $D>1$, we can bound Equation˜59 by $4(\theta^{*})^{-1}(D-1)^{-1}D^{-\frac{1}{2}}a_{0}^{-1}b_{0}^{-D+1}$.
@@
Consider the equation Equation˜38 with $\theta^{*}\geq 0$ ( $a$ similar result holds for $\theta^{*}\leq 0$ ). Suppose $\theta^{*}>8\kappa$. Suppose for some $t_{0}\geq 0$ such that
@@
For all $t\geq t_{0}$, we have $\frac{1}{4}\theta^{*}\leq$ $\theta(t)$ by Lemma˜C.4 (Item 1). Consequently, Equation˜54 implies that
@@
Furthermore, by Equation˜52, we have
@@
Consider the equation Equation˜38 with $\theta^{*}\geq 0$ (a similar result holds for $\theta^{*}\leq 0$ ).
@@
Consider the equation Equation˜38. Denote $r^{\prime}=\min\{a_{0},b_{0}/D\},R^{\prime}=\max\{a_{0},b_{0}/D\}$. W.L.O.G., We assume that $\theta^{*}\geq 0$. Define $T_{1}=\inf\{t:|\beta(t)|>r^{\prime}\}$, and $T_{2}=\inf\{t:|\beta(t)|>R^{\prime}\}$. If $D\geq 1$, and $t$ satisfies the following:
@@
From Equation˜38, we have
@@
Consider $t\leq T_{1}$. We use Equation˜51 to get $a(t)\leq\sqrt{2}a_{0}$ and by Equation˜50, $b^{2}(t)-D\beta^{2}(t)\equiv b_{0}^{2}$. Consequently, we have $b^{2}(t)\leq(1+\frac{1}{D})b_{0}^{2}$, and thus
@@
which implies Equation˜64 by using the fact that $|\theta|=|ab^{D}\beta|$. Furthermore, Equation˜65 implies that
@@
Consider $t\in(T_{1},T_{2})$. We have $a(t)\leq\sqrt{2}|\beta(t)|$ and $b(t)^{2}\leq(1+\frac{1}{D})b_{0}^{2}$. Consequently, Equation˜38 implies that
@@ C.5 Auxiliary Lemma
By ˜5.1, the noise $\xi_{k}$ is sub-Gaussian with variance proxy $C_{\text{proxy}}/n$. Therefore, $\mathbf{P}\left(\left|\xi_{k}\right|\geq s\right)\leq 2\exp(-(2C_{\text{proxy}})^{-1}ns^{2})$.
@@
We combined the Equation˜70 and Equation˜72, and we derive the results.
@@ D.1 Proportional Asymptotics
Several studies analyze Principal Component Regression (PCR) in the proportional asymptotic setting where the dimension $p$ and sample size $n$ grow with $p/n\!\to\!\gamma$. In this regime, Xu and Hsu [46] study the limiting risk of PCR with Gaussian designs with diagonal covariance. They assume polynomially decaying eigenvalues or a convergent empirical spectum, together with an isotropic prior, and they reveal a “double-descent” risk curve. In a related vein, Wu and Xu [45] extend the analysis to a general covariance matrix $\Sigma_{x}$ and an anisotropic prior satisfying $\mathbb{E}\beta_{*}\beta_{*}^{\!\top}\!=\!\Sigma_{\beta}$. They also derive an exact risk expression and demonstrate how “misalignment” between $\Sigma_{x}$ and $\Sigma_{\beta}$ affects risk; here “alignment” refers to concordance between the orderings of their eigenvalues. Both studies assume knowledge of the eigenvectors of the population covariance matrix $\Sigma_{x}$ to construct the oracle PCR. Gedon et al. [17] analyze the limiting risk of PCR under a latent factor model and explore the effect of distribution shift. Green and Romanov [20] derive the exact limits of estimation risk, in-sample prediction risk, and out-of-sample prediction risk of PCR under the assumption that both the empirical spectrum distribution and the distribution of the true signal’s mass over the eigenspaces of $\Sigma_{x}$ converges weakly.
@@ D.2 Non-Asymptotic Analysis
Large Language Models Statement
Large language models were used to improve the clarity of the manuscript and to facilitate experiments.

### v3

@@ C.2 Generalized signal results by dynamic equation analysis
For the choice $b_{0}=cD^{\frac{D+1}{D+2}}\varepsilon^{\frac{1}{D+2}}$, the second term on the right-hand side of Equation˜49 is dominated by the right-hand side of Equation˜50, and we can choose $c$ small enough so that the summation of the two terms is bounded by $b_{0}^{-D}\varepsilon^{-1}\ln n$. This justifies our choice of $t(\varepsilon)$.
@@
Since the events in Lemma C.9 hold, we have $|\xi_{j}|\leq 2(C_{\text{proxy}})^{-1/2}\sqrt{\frac{\ln\tilde{j}+\ln n}{n}}$, where $\tilde{j}=\frac{\tilde{d}}{\lambda_{j}}$. Since $j\in R$, we have $|\theta_{j}^{*}|\leq\tilde{\sigma}\leq\sqrt{c^{\prime}}\varepsilon$. Since $b_{0}^{D}t(\varepsilon)=\varepsilon^{-1}\ln n$, we can check that $t$ is no more than the hitting time $T_{2}$ defined by Equation˜68 in Lemma˜C.8 as follows.
@@ C.3 Conservation quantity
We omit the subscript $j$ in the following two sections C.3 and C.4 because all the proofs are similar for $j=1,2,\cdots,d$. By Equation˜39, it is easy to see that
@@ C.4 Multi-layer dynamic
Consider the equation Equation˜39. Suppose $y>0$.
@@
Items 1 and 2 are directly implied from Equation˜39. Item 3 is implied by Item 2.
To prove Item 4, consider Equation˜53, from which we have
@@
Consider the equation Equation˜39. Suppose $\theta^{*}\geq 8\kappa$ (similar results hold for $\theta^{*}\leq-8\kappa$ by symmetry). Suppose $t_{0}\geq 0$ such that $0\leq\theta\left(t_{0}\right)<\frac{1}{4}\theta^{*}$. Define
@@
From Equation˜39, we have
@@
Case 1: If $a_{0}<b_{0}/\sqrt{D}$, $r=a_{0}$ and $R=b_{0}/\sqrt{D}$. Note $a\geq\beta$ in Equation˜52. We use Equation˜57 to get
@@
Case 2: If $a_{0}>b_{0}/\sqrt{D}$, $R=a_{0}$ and $r=b_{0}/\sqrt{D}$. We use $b\geq\sqrt{D}\beta$ in Equation˜52 together with $a>a_{0}$, $b>\sqrt{D}|\beta|$ and Equation˜57 to get that
@@
Stage 3: If $T^{\textrm{sig}}\leq T^{\textrm{pos},2}$, then we can use the for $T^{\textrm{sig}}$ in Stage 2 as a bound for $T^{\textrm{sig}}$. Now, we consider the case $T^{\textrm{pos},2}<T^{\textrm{sig}}$. We combine Equation˜52 with $a>|\beta|$, $b>\sqrt{D}|\beta|$, and Equation˜57 to get
@@
We now bound $T^{\mathrm{\textrm{sig}}}$ using the summation of Equation˜58, Equation˜61, and Equation˜59 if $a_{0}<b_{0}/\sqrt{D}$, or the summation of Equation˜58, Equation˜61, and Equation˜60 if $a_{0}>b_{0}/\sqrt{D}$.
If $a_{0}<b_{0}/\sqrt{D}$, we can bound the right hand sides of Equation˜58 and Equation˜61 by $8(\theta^{*})^{-1}b_{0}^{-D}$ and $4(\theta^{*})^{-1}b_{0}^{-D}$ respectively.
If $a_{0}>b_{0}/\sqrt{D}$, we can bound the right hand sides of Equation˜58 and Equation˜61 by $8(\theta^{*})^{-1}D^{-\frac{1}{2}}a_{0}^{-1}b_{0}^{-D+1}$ and $4(\theta^{*})^{-1}D^{-\frac{1}{2}}a_{0}^{-1}b_{0}^{-D+1}$ respectively. Furthermore, if $D>1$, we can bound Equation˜60 by $4(\theta^{*})^{-1}(D-1)^{-1}D^{-\frac{1}{2}}a_{0}^{-1}b_{0}^{-D+1}$.
@@
Consider the equation Equation˜39 with $\theta^{*}\geq 0$ ( $a$ similar result holds for $\theta^{*}\leq 0$ ). Suppose $\theta^{*}>8\kappa$. Suppose for some $t_{0}\geq 0$ such that
@@
For all $t\geq t_{0}$, we have $\frac{1}{4}\theta^{*}\leq$ $\theta(t)$ by Lemma˜C.4 (Item 1). Consequently, Equation˜55 implies that
@@
Furthermore, by Equation˜53, we have
@@
Consider the equation Equation˜39 with $\theta^{*}\geq 0$ (a similar result holds for $\theta^{*}\leq 0$ ).
@@
Consider the equation Equation˜39. Denote $r^{\prime}=\min\{a_{0},b_{0}/D\},R^{\prime}=\max\{a_{0},b_{0}/D\}$. W.L.O.G., We assume that $\theta^{*}\geq 0$. Define $T_{1}=\inf\{t:|\beta(t)|>r^{\prime}\}$, and $T_{2}=\inf\{t:|\beta(t)|>R^{\prime}\}$. If $D\geq 1$, and $t$ satisfies the following:
@@
From Equation˜39, we have
@@
Consider $t\leq T_{1}$. We use Equation˜52 to get $a(t)\leq\sqrt{2}a_{0}$ and by Equation˜51, $b^{2}(t)-D\beta^{2}(t)\equiv b_{0}^{2}$. Consequently, we have $b^{2}(t)\leq(1+\frac{1}{D})b_{0}^{2}$, and thus
@@
which implies Equation˜65 by using the fact that $|\theta|=|ab^{D}\beta|$. Furthermore, Equation˜66 implies that
@@
Consider $t\in(T_{1},T_{2})$. We have $a(t)\leq\sqrt{2}|\beta(t)|$ and $b(t)^{2}\leq(1+\frac{1}{D})b_{0}^{2}$. Consequently, Equation˜39 implies that
@@ C.5 Auxiliary Lemma
By ˜6.1, the noise $\xi_{k}$ is sub-Gaussian with variance proxy $C_{\text{proxy}}/n$. Therefore, $\mathbf{P}\left(\left|\xi_{k}\right|\geq s\right)\leq 2\exp(-(2C_{\text{proxy}})^{-1}ns^{2})$.
@@
We combined the Equation˜71 and Equation˜73, and we derive the results.
@@ D.1 Proportional Asymptotics
Several studies analyze Principal Component Regression (PCR) in the proportional asymptotic setting where the dimension $p$ and sample size $n$ grow with $p/n\!\to\!\gamma$. In this regime, Xu and Hsu [46] study the limiting risk of PCR with Gaussian designs with diagonal covariance. They assume polynomially decaying eigenvalues or a convergent empirical spectrum, together with an isotropic prior, and they reveal a “double-descent” risk curve. In a related vein, Wu and Xu [45] extend the analysis to a general covariance matrix $\Sigma_{x}$ and an anisotropic prior satisfying $\mathbb{E}\beta_{*}\beta_{*}^{\!\top}\!=\!\Sigma_{\beta}$. They also derive an exact risk expression and demonstrate how “misalignment” between $\Sigma_{x}$ and $\Sigma_{\beta}$ affects risk; here “alignment” refers to concordance between the orderings of their eigenvalues. Both studies assume knowledge of the eigenvectors of the population covariance matrix $\Sigma_{x}$ to construct the oracle PCR. Gedon et al. [17] analyze the limiting risk of PCR under a latent factor model and explore the effect of distribution shift. Green and Romanov [20] derive the exact limits of estimation risk, in-sample prediction risk, and out-of-sample prediction risk of PCR under the assumption that both the empirical spectrum distribution and the distribution of the true signal’s mass over the eigenspaces of $\Sigma_{x}$ converges weakly.